OpenStudy (anonymous):

Is there a difference between -7² and (-7)²

OpenStudy (driftracer305):

yes ......... with the first one it can be -49........with second one it is just 49

OpenStudy (anonymous):

\[-7^2 = -49, ~~~ (-7)^2 = (-7)(-7) = 49 \]

OpenStudy (anonymous):

Brackets matter :)

OpenStudy (jdoe0001):

\(\bf -7^2\to -1\cdot 7^2\to -1\cdot 7\cdot 7\to -1\cdot 49\to {\color{brown}{ -49}} \\ \quad \\ \quad \\ (-7)^2\to (-7)\cdot (-7)\to {\color{brown}{ +49}}\)